# Leveraging data-driven weather models for improving numerical weather prediction skill through large-scale spectral nudging

### v1

@@ Abstract
Operational meteorological forecasting has long relied on physics-based numerical weather prediction (NWP) models. Recently, this landscape has been disrupted by the advent of data-driven artificial intelligence (AI)-based weather models, which offer tremendous computational performance and competitive forecasting skill. However, data-driven models for medium-range forecasting generally suffer from major limitations, including low effective resolution and a narrow range of predicted variables. This study illustrates the relative strengths and weaknesses of these competing paradigms using the GEM (Global Environmental Multiscale) and GraphCast models to represent physics-based and AI-based approaches, respectively. By analyzing global predictions from these two models against observations and analyses in both physical and spectral spaces, this study demonstrates that GraphCast-predicted large scales outperform GEM, particularly for longer lead times. Building on this insight, a hybrid NWP-AI system is proposed, wherein GEM-predicted large-scale state variables are spectrally nudged toward GraphCast predictions, while allowing GEM to freely generate fine-scale details critical for weather extremes. Results indicate that this hybrid approach is capable of leveraging the strengths of GraphCast to enhance the prediction skill of the GEM model. Importantly, trajectories of tropical cyclones are predicted with enhanced accuracy without significant changes in intensity. Furthermore, this new hybrid system ensures that meteorologists have access to a complete set of forecast variables, including those relevant for high-impact weather events.
@@ 1 Introduction
State-of-the-art numerical weather prediction (NWP) models generally include some form of a dynamical core that solves the nonlinear partial differential equations governing the resolved-scale dynamical and thermodynamical processes within the atmosphere. Solutions from the dynamical core are subsequently combined with tendencies from parameterized physics schemes, representing the unresolved and subgrid-scale processes that are not explicitly accounted for by the dynamical equations. Although statistical approaches have been explored in the past — particularly for downscaling purposes (Yu et al., 2006; Busuioc et al., 2008; Cheng et al., 2014) — physics-based NWP models have long been the foundational approach for operational meteorological forecasting. The recent emergence of data-driven models inspired by artificial intelligence (AI) has, however, started to seriously challenge this well established paradigm.
@@
Keisler (2022) presented an important breakthrough demonstrating considerable potential for weather forecasting with data-driven models. This was quickly followed by Pangu-Weather from Huawei (Bi et al., 2022, 2023), GraphCast from Google-DeepMind (Lam et al., 2022, 2023) and several other models (e.g., Pathak et al. 2022; Chen et al. 2023a, b). In general, these models rely on some form of deep neural network architecture. Like any other application of AI, data-driven weather models necessitate a substantial volume of high quality training data. As a result, all currently available AI-based deterministic global weather simulators are trained on the ERA5 reanalyses (Hersbach et al., 2020) from the European Centre for Medium-Range Weather Forecasts (ECMWF), which is undoubtedly the most comprehensive resource available. The neural network weights within these models are specifically trained to make predictions that closely emulate ERA5.
@@
One important aspect of these AI-based weather emulators that set them apart from the classical NWP models is their computational performance. They can speed up the production of multi-day forecasts by orders of magnitude while using a fraction of the computational resources. Furthermore, many of these models are not only comparable to the NWP models with respect to their prediction accuracy, in terms of a number of the performance metrics, they can even surpass the skill of the IFS (Integrated Forecasting System; Lang et al. 2023) from ECMWF, which is widely accepted as the benchmark within the NWP community (Bi et al., 2023; Lam et al., 2023; Chen et al., 2023a).
@@
Despite their advantages, AI-based weather emulators also have limitations. A major weakness of most of the currently available data-driven models is considerable fine-scale smoothing (Bi et al., 2022; Lam et al., 2022). Notably, data-driven models based on generative AI, e.g., diffusion-based GenCast model from Google-DeepMind (Price et al., 2024), can address the fine-scale smoothing problem to some extent. However, as of now, such approaches generally require considerable increase in computational cost, thereby diminishing one of the principal advantages of the AI models.
@@
Improving the nominal horizontal resolution of the AI model inferences (i.e., forecasts) is also challenging. The general construction of these models requires higher-resolution training data to enable these models to make inferences at higher resolutions. Since the ERA5 reanalysis is only available at 0.25°$\degree$ resolution, this restriction means that current global AI models cannot be trained for kilometer-scale forecasting. Another common limitation of the current generation of AI models is the limited range of predicted variables. At present, these models can only be trained to emulate fields that are present in the analyses dataset, although training to predict non-analyzed variables (e.g., precipitation in ERA5) would generally lead to suboptimal inference. Furthermore, adding the full suite of physics-related prognostic variables required for operational NWP, such as categorization of cloud and precipitation types, would be cost-prohibitive with current AI model architectures. Time resolution of inferences is another constraint for most of the current AI models, which make predictions at 6-hour intervals, far coarser than physics-based NWP models. Pangu-Weather (Bi et al., 2022) provides separate models for 1-, 3-, 6-, and 24-hour predictions, but the models with shorter lead times suffer from larger error accumulation. In summary, despite their strengths, the aforementioned limitations make it implausible for the AI models to replace the classical NWP approach in operational weather forecasting.
The GEM (Global Environmental Multiscale) model (Girard et al., 2014), which is employed by Environment and Climate Change Canada (ECCC) for operational weather forecasting, relies entirely on the physics-based NWP approach. The principal objective of this study is to determine whether inferences from a state-of-the-art AI-based weather emulator can be leveraged through large-scale spectral nudging to enhance the predictability of the GEM model, particularly for global weather forecasting. Spectral nudging can be seen as an indirect suboptimal data assimilation method (von Storch et al., 2000). Although the concept of spectral nudging was first proposed to control spatial computational modes in a limited-area model (LAM; Waldron et al. 1996), it is primarily used to improve dynamical downscaling of coarse-resolution atmospheric model simulations over high-resolution limited area domains (Leduc and Laprise, 2009; Husain et al., 2014). Spectral nudging has also been applied for global atmospheric hindcasting (Schubert-Frisius et al., 2017).
@@
The ultimate goal of this study is to develop a hybrid NWP-AI system that combines the strengths of both approaches while addressing the inherent limitations of the AI-based models. The AI model chosen for this study is the GraphCast model from Google-DeepMind. Based on internal testing at ECCC, among the different freely available open-source AI-based models, GraphCast demonstrates the best accuracy with respect to most of the metrics.
@@
Spectral nudging has so far been unfeasible for real-time forecasting applications due to the lack of availability of timely and accurate reference data. The emergence of AI models like GraphCast has unlocked the potential for applying spectral nudging in real-time operational weather forecasting. In this regard, application of spectral nudging for enhancing real-time NWP represents an important novelty proposed by this study. Although this paper illustrates the hybridization concept with a specific combination of the AI and NWP models, the proposed strategy is expected to be broadly applicable to any other combination as long as the AI model offers some added value over the NWP model.
This article is organized as follows. Sections 2 and 3 provide some important background information regarding the GEM and GraphCast models, respectively. Details pertaining to the relative performances of these models are documented in section 4. Section 5 covers the concept of spectral nudging, its implementation in GEM as well as the optimal nudging configuration used in this study. Detailed evaluations of the spectral nudging-based hybrid NWP-AI system are presented in 6. Finally, section 7 summarizes the conclusions of this study and outlines potential directions for future work.
@@ 2.1 The GEM Model
The dynamical core of the GEM model solves the elastic Euler system of equations. These equations are first transformed from regular height coordinate to some form of a terrain-following coordinate (TFC) in the vertical, denoted by $\zeta$ in the model. All current operational GEM-based NWP systems employ a log-hydrostatic-pressure-type TFC (Girard et al., 2014; Husain et al., 2021).
The nonhydrostatic version of the dynamical core involves eight prognostic variables: horizontal velocity components ($u$ and $v$), vertical velocity ($w$), virtual temperature ($T_{v}$), geopotential ($\phi$), vertical motion with respect to the TFC ($\dot{\zeta}$), hydrostatic surface pressure ($\pi_{s}$), and nonhydrostatic pressure perturbation ($q=\ln\frac{p}{\pi}$, where $p$ and $\pi$ denote the complete and hydrostatic pressure, respectively).
The advection terms in the model are treated with semi-Lagrangian approach while a two-time-level iterative implicit Crank-Nicolson type scheme is employed for temporal discretization. Spatial discretization relies on staggered grids: an Arakawa C grid in the horizontal and a Charney-Phillips grid in the vertical. Further details in this regard can be found in the existing literature (Girard et al., 2014; Husain and Girard, 2017). Global simulations in GEM utilize a Yin-Yang grid system, where two orthogonal and overlapping LAM grids are combined to construct the global grid (Qaddouri and Lee, 2011).
Solutions from the GEM dynamical core are combined with tendencies from parameterized physics schemes that represent unresolved (or partially resolved) physical processes, such as radiation, convection, boundary layer turbulence, surface-atmosphere interactions, condensation, etc. The parameterization schemes are executed in a sequential manner. All operational GEM-based NWP systems rely on a splitting-type mechanism for coupling the parameterized physics tendencies with the resolved dynamics. In this approach, physics-induced tendencies are applied as grid-point adjustments to the dynamics solution to obtain the complete solution for a model time step (Husain et al., 2019).
@@ 2.2 Global Deterministic Prediction System (GDPS)
The version of the GDPS used in this study is based on its current operational configuration. It has a horizontal grid spacing of approximately 15 km. In the vertical, GDPS employs 84 prognostic vertical levels with the topmost level at approximately 0.1 hPa. The first momentum and thermodynamical levels are approximately 20 m and 10 m away from the surface, respectively. It uses a time-step length of 450 s.
In terms of the underlying physics parameterizations, GDPS has undergone a major modernization in the recent past (McTaggart-Cowan et al., 2019). This includes updated versions of schemes for radiative transfer, blocking, orography gravity waves and non-orographic gravity waves.
Although operational GDPS is coupled with the NEMO ocean model (Smith et al., 2018), all the experiments for this study were conducted with atmosphere-only configurations. As the impact of coupling is usually modest for 10-day forecasts (McTaggart-Cowan et al., 2019), conclusions drawn from this study are expected to generally hold in the presence of atmosphere-ocean coupling.
@@ 3.1 The GraphCast Model
The GraphCast (GC) model from Google-DeepMind is an AI-based weather emulator which has been trained to emulate ECMWF’s ERA5 reanalyses. It requires the state of the atmosphere at times $t$ and $t-\Delta t_{GC}$ to produce a forecast valid at $t+\Delta t_{GC}$, where $\Delta t_{GC}$=6 hr. Although the model can be employed auto-regressively $N$ times to generate predictions valid at $t+N\Delta t_{GC}$, it is trained to minimize error for up to 12 recursive forecast steps (from 6 hr to 3 days) by considering ERA5 as the ground truth. The objective of such a training is to produce forecasts with reasonable skill for up to day 10.
@@
The atmospheric state predicted by GC is represented by 6 variables defined on multiple pressure levels (temperature, $u-v$ components of wind, geopotential, specific humidity, and vertical wind speed) and 5 surface variables (2-m temperature, 10-m wind components, mean sea-level pressure, and total precipitation). In addition to the predicted variables, the input system of GC requires time-dependent forcing in the form of total incident solar radiation at the top of the atmosphere along with information regarding the local time of the day and progression of year. These time-dependent forcing terms are computed analytically. Furthermore, the input system requires some stationary input features like the land-sea mask, surface geopotential as well as information on the local longitude and latitude of its input grid.
Pre-trained weights of GC (learned features of the GNNs) are available from Google-DeepMind for two different sets of pressure levels. The version with higher vertical resolution has 37 pressure levels with the top at 1 hPa, while the coarser-resolution version has only 13 levels with the top at 50 hPa (Lam et al., 2022). It is important to note that, in addition to training on ERA5, the 13-level GC was also fine-tuned by DeepMind on initial conditions for ECMWF’s operational IFS system, from 2016 to 2021.
@@ 3.2 GraphCast-based Global Prediction System
Based on the evaluation of GC predictions with the GDPS analyses as input, the 13-level version of GC (GC-13L) outperforms the 37-level version (GC-37L) for most variables, at most levels and over most regions of the world. This is presumably due to the fact that stratospheric specific humidity in the GDPS analyses generally differs from ERA5 by an order of magnitude creating large imbalances when 37-level GC (with multiple levels in the stratosphere) is initialized with the GDPS analyses. The previously mentioned additional fine-tuning of GC-13L by DeepMind could also have contributed to its superior performance. This study is, therefore, entirely based on the GC-13L and, henceforth, GC is used to refer to GC-13L.
@@ 4 Relative performances of the GDPS and GC forecasts
4.1 Experimental setup
Ten-day forecasts from GC and GDPS are initialized using the GDPS analysis every 36 h, leading to 60 cases for boreal winter 2022 (from 01 December 2021 at $0$ UTC to 27 February 2022 at 1200 UTC) and 62 cases for boreal summer 2022 (from 01 June 2022 at $0$ UTC to 31 August 2022 at 1200 UTC). This version of GDPS, which includes data assimilation and forecasting components, became operational in June 2024 (ECCC, 2024). Henceforth, the GDPS experiment based on this configuration will be referred to as GDPS-CTL.
@@ 4.2 Verification with global power spectra
Although GC employs a 0.25°$\degree$ latitude-longitude resolution grid, its predictions are subject to a smoothing effect, which increases with lead time. This is generally assumed to be associated with learning to minimize the mean square error (MSE) (e.g., Keisler (2022), Lam et al. (2023)). The smoothing effect is well indicated by departures of 120-h GC forecasts’ global 250-hPa kinetic energy and 500-hPa temperature variance spectra from those of GDPS-CTL forecasts and the analyses (Fig. 1). GraphCast’s spectra are steeper, leading to a notable variance deficiency at wavenumbers as low as 30, corresponding to lower synoptic and sub-synoptic scales ($<$1500 km). However, both GDPS-CTL forecasts and analyses much more closely follow the $k^{-3}$-law, pertaining to the 3-dimensional quasi-geostrophic flow (Charney, 1971; Skamarock et al., 2014), which indicates that GC slopes are not realistic. Hence, the effective resolution of GC, defined as the highest wave number at which the atmospheric variability is still adequately represented, is considerably lower than its nominal resolution. Conversely, the effective resolution of the classical NWP models is typically about 6$\sim$8 times their nominal resolution, i.e., grid spacing (Skamarock, 2004). With the current operational configuration, the nominal resolution for GDPS-CTL is 0.1375°$\degree$, leading to an effective resolution of around 90-100 km.
Figure 1: Global (a) kinetic energy power spectrum of 250-hPa horizontal wind and (b) variance spectrum of 500-hPa temperature after 120 h of integration for the GDPS analyses (black), GDPS-CTL (red) and GC (red), averaged over 60 cases of boreal winter 2022. The dashed line indicates the -3 slope.
@@
$\gamma=\left(\frac{\overline{\sigma_{x^{\prime}x^{\prime}}}}{\overline{\sigma_{y^{\prime}y^{\prime}}}}\right)^{1/2},$ (1)
where $\sigma_{x^{\prime}x^{\prime}}$ and $\sigma_{y^{\prime}y^{\prime}}$ are the power spectra of the forecast’ and analysis’ transient-eddy components, respectively, and the overbar denotes averaging over the cases for a given season. The forecast transient eddies are defined as $x^{\prime}=x-\overline{x}$, and similarly for the analysis. Figure 2 shows the evolution of $\gamma$ with lead time for GC and GDPS-CTL forecasts, against the GDPS analysis. While $\gamma$ remains close to 1 for all lead times for GDPS-CTL, indicating that its effective resolution does not change during the integration, for GC it rapidly decreases, leading to an effective resolution as low as 1000 km for 24-h forecasts. The smoothing effect further reduces the effective resolution to 2500$\sim$2700 km before it saturates around forecast days 3–5. It is worth noting here that GC still resolves scales smaller than these limits, but only partly ($10<\gamma<$ 1). While designing any strategy to leverage GC predictions to improve GEM, it would be prudent to not force the GEM model to follow GC for these variance-deficient scales in order to not adversely affect GEM’s effective resolution. This will be revisited in Section 5, where the concept of spectral nudging and its implementation in GEM are discussed.
Figure 2: Spectral (a, b) amplitude ratio $\gamma$ and (c, d) coherence $\rho$ for the global (a, c) kinetic energy spectrum of 500-hPa horizontal wind transient-eddy deviations and (b, d) variance spectrum of 500-hPa temperature transient-eddy deviations, or GDPS-CTL (blue) and for GC (red), after (dark to light) 24, 72, 120 and 240 h of integration. Spectra are averaged over 60 cases of boreal winter 2022.
@@
$\rho=\frac{\overline{\sigma_{x^{\prime}y^{\prime}}}}{\left(\overline{\sigma_{x^{\prime}x^{\prime}}}\;\overline{\sigma_{y^{\prime}y^{\prime}}}\right)^{1/2}}\;,$ (2)
where $\sigma_{x^{\prime}y^{\prime}}$ is the transient-eddy co-spectrum between the GDPS-CTL or GC forecasts and the GDPS analyses. Since the same GEM model is used to produce both the GDPS forecasts and backgrounds for the analysis steps, they may have common systematic errors, especially at shorter lead times. This may unfairly penalize GC, which is trained on ECMWF-ERA5 reanalyses. Despite that, Fig. 2 clearly indicates an improved spectral coherence with the GDPS analyses for a large range of scales – from synoptic to planetary, especially at longer lead times. This suggests that forcing GEM toward GC has potential for improving the predictive skill of its transient-eddy anomalies.

### v2

@@ Abstract
Operational meteorological forecasting has long relied on physics-based numerical weather prediction (NWP) models. Recently, this landscape is facing disruption by the advent of data-driven artificial intelligence (AI)-based weather models, which offer tremendous computational performance and competitive forecasting skill. However, data-driven models for medium-range forecasting generally suffer from major limitations, including low effective resolution and a narrow range of predicted variables. This study illustrates the relative strengths and weaknesses of these competing paradigms using the GEM (Global Environmental Multiscale) and GraphCast models to represent physics-based and AI-based approaches, respectively. By analyzing global predictions from these two models against observations and analyses in both physical and spectral spaces, this study demonstrates that GraphCast-predicted large scales outperform GEM, particularly for longer lead times. Building on this insight, a hybrid NWP-AI system is proposed, wherein GEM-predicted large-scale state variables are spectrally nudged toward GraphCast predictions, while allowing GEM to freely generate fine-scale details critical for weather extremes. Results indicate that this hybrid approach is capable of leveraging the strengths of GraphCast to enhance the prediction skill of the GEM model. Importantly, trajectories of tropical cyclones are predicted with enhanced accuracy without significant changes in intensity. Furthermore, this new hybrid system ensures that meteorologists have access to a complete set of forecast variables, including those relevant for high-impact weather events.
@@ 1 Introduction
State-of-the-art numerical weather prediction (NWP) models generally include some form of a dynamical core that solves the nonlinear partial differential equations governing the resolved-scale dynamical and thermodynamical processes within the atmosphere. Solutions from the dynamical core are subsequently combined with tendencies from parameterized physics schemes, representing diabatic, frictional and subgrid-scale processes that are not explicitly accounted for by the dynamical equations. Although statistical approaches have been explored in the past — particularly for downscaling purposes (Yu et al., 2006; Busuioc et al., 2008; Cheng et al., 2014) — physics-based NWP models have long been the foundational approach for operational meteorological forecasting. The recent emergence of data-driven models inspired by artificial intelligence (AI) has, however, started to seriously challenge this well established paradigm.
@@
Keisler (2022) presented an important breakthrough demonstrating considerable potential for weather forecasting with data-driven models. This was quickly followed by Pangu-Weather from Huawei (Bi et al., 2022, 2023), GraphCast from Google DeepMind (Lam et al., 2022, 2023) and several other models (e.g., Pathak et al. 2022; Chen et al. 2023a, b). In general, these models rely on some form of deep neural network architecture. Like any other application of AI, data-driven weather models necessitate a substantial volume of high quality training data. As a result, all currently available AI-based deterministic global weather simulators are trained on the ERA5 reanalyses (Hersbach et al., 2020) from the European Centre for Medium-Range Weather Forecasts (ECMWF), which is undoubtedly the most comprehensive resource available. The neural network weights within these models are specifically trained to make predictions that closely emulate ERA5.
@@
One important aspect of these AI-based weather emulators that set them apart from the classical NWP models is their computational performance. They can speed up the production of multi-day forecasts by orders of magnitude while using a fraction of the computational resources. Furthermore, many of these models are not only comparable to the NWP models with respect to their prediction accuracy, in terms of a number of the performance metrics, they can even surpass the skill of the Integrated Forecasting System (IFS; Lang et al. 2023) from ECMWF, which is widely accepted as the benchmark within the NWP community (Bi et al., 2023; Lam et al., 2023; Chen et al., 2023a).
@@
Despite their advantages, AI-based weather emulators also have limitations. A major weakness of most of the currently available data-driven models is considerable fine-scale smoothing (Bi et al., 2022; Lam et al., 2022). Data-driven models based on generative AI, e.g., diffusion-based GenCast model from Google DeepMind (Price et al., 2024), can address the fine-scale smoothing problem to some extent. However, as of now, such approaches generally require considerable increase in computational cost, thereby diminishing one of the principal advantages of the AI models.
@@
Improving the nominal horizontal resolution of the AI model inferences (i.e., forecasts) is also challenging. The general construction of these models requires higher-resolution training data to enable these models to make inferences at higher resolutions. Since the ERA5 reanalysis is only available at 0.25°$\degree$ resolution, this restriction means that current global AI models cannot be trained for kilometer-scale forecasting. Another common limitation of the current generation of AI models is the limited range of predicted variables. At present, these models can only be trained to emulate fields that are present in the analyses dataset. However, training to predict non-analyzed variables present in the analyses (e.g., precipitation in ERA5) would generally lead to suboptimal inference. Furthermore, adding the full suite of physics-related prognostic variables required for operational NWP, such as categorization of cloud and precipitation types, would be cost-prohibitive with current AI model architectures. Time resolution of inferences is another constraint for most of the current AI models, which make predictions at 6-hour intervals, far coarser than physics-based NWP models. Pangu-Weather (Bi et al., 2022) provides separate models for 1-, 3-, 6-, and 24-hour predictions, but the models with shorter forecasting steps suffer from larger error accumulation. In summary, despite their strengths, the aforementioned limitations make it implausible for the AI models to replace the classical NWP approach in operational weather forecasting.
@@
The ultimate goal of this study is to develop a hybrid NWP-AI system for real-time forecasting applications that combines the strengths of the physics-based NWP approach with those of the AI-based approach while addressing the inherent limitations of the AI models. The physics-based model chosen for this study is the GEM (Global Environmental Multiscale) model (Girard et al., 2014), which is employed by Environment and Climate Change Canada (ECCC) for operational weather forecasting. The AI model in this study is represented by GraphCast from Google DeepMind. Based on internal testing at ECCC, among the different freely available open-source AI models, GraphCast demonstrates the best accuracy with respect to most of the metrics.
The study aims to determine whether large-scale spectral nudging of GEM predictions toward inferences from a state-of-the-art AI-based weather emulator, such as GraphCast, can improve GEM’s predictability and offer a feasible approach for developing a suitable hybrid NWP-AI system. Spectral nudging can be seen as an indirect suboptimal data assimilation method (von Storch et al., 2000). Although the concept of spectral nudging was first proposed to control spatial computational modes in a limited-area model (LAM; Waldron et al. 1996), it is primarily used to improve dynamical downscaling of coarse-resolution atmospheric model simulations over high-resolution limited area domains (Leduc and Laprise, 2009; Husain et al., 2014). Spectral nudging has also been explored for global atmospheric hindcasting (Schubert-Frisius et al., 2017).
Spectral nudging has so far been unfeasible for real-time forecasting applications due to the absence of timely and accurate reference data. The emergence of AI models, like GraphCast, has unlocked the potential for applying spectral nudging in real-time operational weather forecasting. In this context, application of spectral nudging for enhancing real-time NWP represents an important novelty proposed by this study. Although the hybridization concept is illustrated with a specific combination of AI and NWP models, the proposed strategy is expected to be broadly applicable to any other combination, provided that the AI model offers added value over the NWP model.
@@
This article is organized as follows. Sections 2 and 3 provide some important background information regarding the GEM and GraphCast models, respectively. Details pertaining to the relative performances of these models are documented in section 4. Section 5 covers the concept of spectral nudging, its implementation in GEM as well as the optimal nudging configuration used in this study. Detailed evaluations of the spectral nudging-based hybrid NWP-AI system are presented in 6. Finally, section 7 summarizes the main conclusions of this study and outlines potential directions for future work.
@@ 2.1 The GEM Model
The dynamical core of the GEM model solves the elastic Euler system of equations using an iteratively implicit semi-Lagrangian approach (Girard et al., 2014; Husain and Girard, 2017). The equations are first transformed from regular height coordinate to some form of a terrain-following coordinate in the vertical, denoted by $\zeta$ in the model, which for ECCC’s current operational NWP systems, is based on log-hydrostatic-pressure (Husain et al., 2021). The global domain is represented as a pair of overlapping limited-area grids in a Yin-Yang configuration (Qaddouri and Lee, 2011).
Solutions from the GEM dynamical core are combined with tendencies from parameterized physics schemes (McTaggart-Cowan et al., 2019) that represent the diabatic, frictional and subgrid-scale processes. The parameterization schemes are executed in a sequential manner. Sequential coupling (Gross et al., 2018), also referred to as split-type coupling (Husain et al., 2019), is employed to combine the parameterized physics tendencies with the resolved dynamics to obtain the complete solution for a model time step.
@@ 2.2 Global Deterministic Prediction System (GDPS)
The version of the GDPS used in this study is based on its current operational configuration (McTaggart-Cowan et al., 2019; ECCC, 2024). It has a horizontal grid spacing of approximately 15 km along with a time-step length of 450 s. In the vertical, GDPS employs 84 prognostic levels with the top approximately at 0.1 hPa.
Although operational GDPS is coupled with the NEMO ocean model (Smith et al., 2018), all the experiments for this study were conducted with atmosphere-only configurations. This simplification reduces both complexity and computational cost, and in the past, has been found to provide reliable guidance for the behavior of the full operational system.
@@ 3.1 The GraphCast Model
The GraphCast (GC) model from Google DeepMind is based on AI, and like other global AI weather models, has been trained to emulate ECMWF’s ERA5 reanalyses (Lam et al., 2022, 2023). It requires the state of the atmosphere at times $t$ and $t-\Delta t_{GC}$ to produce a forecast valid at $t+\Delta t_{GC}$, where $t$ is the initialization time and $\Delta t_{GC}$=6 hr. Although the model can be employed auto-regressively $N$ times to generate predictions valid at $t+N\Delta t_{GC}$, it is trained to minimize the mean squared error (weighted by vertical levels) for up to 12 recursive forecast steps (from 6 hr to 3 days) by considering ERA5 as the ground truth. The objective of such a training is to produce forecasts with reasonable skill for up to day 10.
@@
The atmospheric state predicted by GC is represented by 6 variables defined on multiple pressure levels (temperature, $u-v$ components of wind, geopotential, specific humidity, and vertical wind speed) and 5 surface variables (2-m temperature, 10-m wind components, mean sea-level pressure, and total precipitation). In addition to the predicted variables, the input system of GC requires time-dependent forcing in the form of total incident solar radiation at the top of the atmosphere along with information regarding the local time of the day and progression of year. These time-dependent forcing terms are computed analytically. Furthermore, the input system requires some stationary input features, such as land-sea mask, surface geopotential as well as information on the local longitude and latitude of its input grid.
Pre-trained weights of GC (learned features of the GNNs) are available from Google DeepMind for two different sets of pressure levels. The version with higher vertical resolution has 37 pressure levels with the top at 1 hPa, while the coarser-resolution version has only 13 levels with the top at 50 hPa (Lam et al., 2022). It is important to note that, in addition to training on ERA5, the 13-level GC was also fine-tuned by DeepMind on initial conditions for ECMWF’s operational IFS system, from 2016 to 2021.
@@ 3.2 GraphCast-based Global Prediction System
Based on the evaluation of GC predictions with the GDPS analyses as input, the 13-level version of GC outperforms the 37-level version for most variables, at most levels and over most regions of the world. Although a detailed analysis of this behavior is beyond the scope of this work, it is believed to be related to systematic differences in stratospheric moisture values between the initializing GDPS analyses and ERA5 reanalyses used to train GC. The previously mentioned additional fine-tuning of GC-13L by DeepMind could also have contributed to its superior performance. This study is therefore, entirely based on the 13-level version of the system and the acronym ‘GC’ is henceforth used to refer to this version.
@@ 4 Relative performances of the GDPS and GC forecasts
Ten-day forecasts from GC and GDPS are initialized using the GDPS analysis every 36 h, leading to 60 cases for boreal winter 2022 (from 01 December 2021 at $0$ UTC to 27 February 2022 at 1200 UTC) and 62 cases for boreal summer 2022 (from 01 June 2022 at $0$ UTC to 31 August 2022 at 1200 UTC). The sequence of forecast integrations initialized from these dates using an atmosphere-only operational GDPS configuration is denoted GDPS-CTL.
Predictions from the two systems, GDPS-CTL and GC, are first assessed for their ability to depict the full range of atmospheric circulations using variance spectra (section 4a). This analysis is aimed at ascertaining effective resolutions of the two models, which will help to fairly interpret their relative predictive skills when comparing against radiosonde observations later in section 4b.
@@ 4.1 Verification with global power spectra
Although GC employs a 0.25°$\degree$ latitude-longitude resolution grid, its predictions are subject to a smoothing effect, which increases with lead time. This is generally assumed to be associated with learning to minimize the mean square error (MSE) (e.g., Keisler (2022), Lam et al. (2023)). The smoothing effect is well indicated by departures of 120-hr GC forecasts’ global 250-hPa kinetic energy and 500-hPa temperature variance spectra from those of GDPS-CTL forecasts and the analyses (Fig. 1). For the fine scales, GraphCast’s spectra are steeper, leading to a notable variance deficiency at wavenumbers as low as 30, corresponding to lower synoptic and sub-synoptic scales ($<$1500 km). More importantly, both GDPS-CTL forecasts and analyses much more closely follow the expected $k^{-3}$ spectral slope pertaining to the 3-dimensional quasi-geostrophic flow (Charney, 1971; Skamarock et al., 2014), which indicates that GC slopes are not realistic. Hence, the effective resolution of GC, defined as the highest wave number at which the atmospheric variability is still adequately represented, is considerably lower than its nominal resolution, i.e., grid spacing. Conversely, the effective resolution of the classical NWP models is typically about 6$\sim$8 times their nominal resolution (Skamarock, 2004). With the current operational configuration, the nominal resolution for GDPS-CTL is 0.1375°$\degree$, leading to an approximate effective resolution of 90$\sim$120 km (McTaggart-Cowan et al., 2024).
Figure 1: Global (a) kinetic energy power spectrum of 250-hPa horizontal wind and (b) variance spectrum of 500-hPa temperature for the GDPS analyses (black) as well as 120-hr lead time of GDPS-CTL (blue) and GC (red), averaged over 60 cases of boreal winter 2022. The dashed line indicates the -3 slope. The 5%–95% confidence interval is presented with light shadings (not discernable for the given logarithmic scale).
@@
$\gamma(n)=\left(\frac{\overline{\sigma_{x^{\prime}x^{\prime}}(n)}}{\overline{\sigma_{y^{\prime}y^{\prime}}(n)}}\right)^{1/2},$ (1)
where $\sigma_{x^{\prime}x^{\prime}}$ and $\sigma_{y^{\prime}y^{\prime}}$ are real-valued non-negative functions of the total spherical wavenumber $n$, representing global spectral variances of the transient-eddy components associated with the forecast and analysis, respectively. The overbar denotes averaging over the cases for a given season. The forecast transient eddies are defined as $x^{\prime}=x-\overline{x}$, and similarly for the analysis. The optimal value of $\gamma$ is 1 for the range of scales that are adequately represented in the reference analysis. In general, for any given length scale (or wavenumber), a value of $\gamma$ larger than 1 would imply that the forecast has more variability than the analysis, whereas a value smaller than 1 would indicate variance deficiency in the forecast.
Figure 2 shows the evolution of $\gamma$ with lead time for GC and GDPS-CTL forecasts against the GDPS analysis. While $\gamma$ remains close to 1 for all lead times for GDPS-CTL, indicating that its effective resolution does not change during the integration, for GC it rapidly decreases, leading to an effective resolution as low as 1000 km for 24-hr forecasts. The smoothing effect further reduces the effective resolution to 2500$\sim$2700 km before it saturates around forecast days 3$\sim$5. It is worth noting here that GC still resolves scales smaller than these limits, but only partly ($\gamma<1$). The impact of this variance-deficiency on optimal usage of GC guidance in a hybrid system will be discussed in more detail in section 5.
Figure 2: Spectral (a, b) amplitude ratio $\gamma$ and (c, d) coherence $\rho$ for the global (a, c) kinetic energy spectrum of 500-hPa horizontal wind transient-eddy deviations and (b, d) variance spectrum of 500-hPa temperature transient-eddy deviations, for GDPS-CTL (blue) and for GC (red), after 24, 72, 120 and 240 hr of integration (darker to lighter shades of color). Spectra are averaged over 60 cases of boreal winter 2022. Light shadings indicate the 5%–95% confidence interval.
@@
$\rho(n)=\frac{\overline{\sigma_{x^{\prime}y^{\prime}}(n)}}{\left(\overline{\sigma_{x^{\prime}x^{\prime}}(n)}\;\overline{\sigma_{y^{\prime}y^{\prime}}(n)}\right)^{1/2}}\;,$ (2)
where $\sigma_{x^{\prime}y^{\prime}}$ is a real-valued function of the total spherical wavenumber $n$, representing the transient-eddy covariance spectrum between the GDPS-CTL (or GC forecasts) and the GDPS analyses. Spectral coherence can be interpreted as a scale-dependent spatial correlation coefficient with an optimal value of 1, which would imply a perfect spatial correlation between forecast and analysis transient-eddy anomalies. Conversely, a value of 0 would imply a total de-correlation. By integrating the individual terms on the right-hand side of Eq. 2 across entire spectrum of wavenumbers, one may obtain a quantity which is analogous to the anomaly correlation, but in space.
Since the same GEM model is used to produce both the GDPS forecasts and backgrounds for the analysis steps, they may have common systematic errors, especially at shorter lead times. Even though this penalizes the ERA5-trained GC, Figs. 2c-d indicate that GC long-lead inferences are superior to the equivalent GDPS predictions over a broad range of synoptic and planetary scales.